OpenStudy (anonymous):

Can someone check 12 questions for me? Please? :) Scientific Notation is the subject.

OpenStudy (anonymous):

What is 9.46 × 10^7 expressed in standard notation? A. 0.000000946 B. 0.0000946 C. 94,600,000 <--------- D. 9,460,000,000 What is 0.000055 expressed in scientific notation? A. 5.5 × 10^–5 <----------- B. 5.5 × 10^–4 C. 5.5 × 10^4 D. 5.5 × 10^5 @mathslover

mathslover (mathslover):

All are right now. Go ahead.

OpenStudy (anonymous):

There are 8.45 × 10^5 registered voters in a certain town. Half of those people are expected to vote in the upcoming election. How many people are expected to vote in the upcoming election? A. 4.225 × 10^2 people B. 4.225 × 10^3 people C. 4.225 × 10^4 people D. 4.225 × 10^5 people <---------- Holly's city has 2.3 × 10^6 residents. Michael’s city has 3.5 × 10^5 fewer residents than Holly's city. How many people live in Michael’s town? A. 1.95 × 10^2 people B. 1.95 × 10^5 people C. 1.95 × 10^6 people <--------------------- D. 1.95 × 10^7 people

mathslover (mathslover):

Yes! First one is right. Second one.. .Hmm! 23 * 10^5 = Holly's city Michael's city = 3.5 * 10^5 Subtract and you get = ?

mathslover (mathslover):

Second one is right too! :-) (was just verifying myself)

OpenStudy (anonymous):

1950000

OpenStudy (anonymous):

A certain kind of passenger jet can travel at a speed of 8.4 × 10^5 meters in 1 hour. At this speed, how far can the jet travel in 5 hours? A. 4.2 × 10^4 meters B. 4.2 × 10^5 meters C. 4.2 × 10^6 meters <------------- D. 8.4 × 10^10 meters In 2007, China had a population of about 1.32 × 10^9. The United States had a population of about 3.01 × 10^8. Using these estimates, how many more people in 2007 lived in China than the United States? A. 1.019 × 10^8 B. 1.019 × 10^9 <---------------- C. 1.69 × 10^8 D. 4.33 × 10^9

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

Both are right Go ahead.

OpenStudy (anonymous):

What is 4,500,000 expressed in scientific notation? A. 4.5 × 10^6 <------ B. 45 × 10^5 C. 450 × 10^8 D. 450 × 10^4 What is 500,000 × 6,000 expressed in scientific notation? A. 3 × 10^6 B. 3 × 10^7 C. 3 × 10^8 D. 3 × 10^9 <------- @mathslover

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

yes... both are right. sorry too many tags :P

OpenStudy (anonymous):

s'ok. :) In the year 2000, the United States government budgeted $1,023,000,000 for the Food and Drug Administration (FDA). What is 1,023,000,000 in scientific notation? A. 1.023 × 10^6 B. 1.023 × 10^7 C. 1.023 × 10^8 D. 1.023 × 10^9 <------- What is 0.00000628 in scientific notation? A. 6.28 • 10^−5 B. 6.28 • 10^−6 <-------- C. 6.28 • 10^5 D. 6.28 • 10^6 @mathslover

mathslover (mathslover):

Corrrrrect!

OpenStudy (anonymous):

There are 8.73 × 10^8 people in the world who speak Mandarin Chinese. There are 3.22 × 10^8 people who speak Spanish. How many more people speak Mandarin Chinese than Spanish? A. 5.51 × 10^4 B. 5.51 × 10^5 C. 5.51 × 10^8 <------ D. 5.51 × 10^16 (8.88 × 104) + (7.48 × 104) = __________ A. 16.36 × 10^3 B. 1.636 × 10^4 <------- C. 1.636 × 10^5 D. 16.36 × 10^8 @mathslover

mathslover (mathslover):

First is right, second is wrong. Check it again.

OpenStudy (anonymous):

A? @mathslover

mathslover (mathslover):

Nope! See, we have : \( 10^4 ( 8.88 + 7.48 ) \) = \(10^4 \times 16.36 \) Now, 16.36 * 10^4 can be written as = ?

OpenStudy (anonymous):

163600

mathslover (mathslover):

I meant to say, in scientific notation. 16.36 * 10^4 = 16.36 * 10^4 \(\times \cfrac{10}{10}\) Now, this becomes : 16.36 / 10 * 10^4 * 10 1.636 * 10^5

OpenStudy (anonymous):

Which = 163600 So... C? @mathslover

mathslover (mathslover):

Yep.. C

mathslover (mathslover):

Game over? :P I think 12 questions are done... !

OpenStudy (anonymous):

Thanks! I'm going to bed, it's 4 am. Night! :)

mathslover (mathslover):

Sure. Good Night. Sweet Dreams! :-)